OpenStudy (anonymous):

Applied Calc- If an object is launched straight up into the air from a starting height of h_{0} feet, then the height of the object after t seconds is approximately h=-16t^2+v_{0}t+h_{0} feet, where v_{0} is the inital velocity of the object. Find the starting height and inital velocity of an object that attains a maximum height of 412 feet five seconds after being launched.

OpenStudy (anonymous):

\[h(t)=-16t^2+v_{0}t+h_{0}\] since max occurs at vertex which is \(-\frac{b}{2a}=\frac{v_0}{32}=5\) we can solve for \(v_0\)

OpenStudy (anonymous):

we get \(v_0=160\) and now you know \[h(5)=412=-16\times 25+160\times 5+h_0\] \[412=-400+800+h_0\] and we can solve for \(h_0\)

OpenStudy (anonymous):

i guess \(h_0=12\)

OpenStudy (anonymous):

didn't really use any calc, did we?

OpenStudy (anonymous):

i completley forgot about the vertex formula, thank you so much